### v1

# DualAnoDiff: Dual-Interrelated Diffusion Model for Few-Shot Anomaly Image Generation

Abstract
The performance of anomaly inspection in industrial manufacturing is constrained by the scarcity of anomaly data. To overcome this challenge, researchers have started employing anomaly generation approaches to augment the anomaly dataset. However, existing anomaly generation methods suffer from limited diversity in the generated anomalies and struggle to achieve a seamless blending of this anomaly with the original image. In this paper, we overcome these challenges from a new perspective, simultaneously generating a pair of the overall image and the corresponding anomaly part. We propose DualAnoDiff, a novel diffusion-based few-shot anomaly image generation model, which can generate diverse and realistic anomaly images by using a dual-interrelated diffusion model, where one of them is employed to generate the whole image while the other one generates the anomaly part. Moreover, we extract background and shape information to mitigate the distortion and blurriness phenomenon in few-shot image generation. Extensive experiments demonstrate the superiority of our proposed model over state-of-the-art methods in terms of both realism and diversity. Overall, our approach significantly improves the performance of downstream anomaly detection tasks, including anomaly detection, anomaly localization, and anomaly classification tasks.
Introduction
Industrial anomaly inspection, i.e., anomaly detection, localization, and classification, plays an important role in industrial manufacture (Duan et al. 2023). However, in real-world industrial production, anomaly samples are scarce. Therefore, the current mainstream anomaly inspection methods are either unsupervised methods which use only normal samples (Roth et al. 2022; Li et al. 2024) or semi-supervised method (Zhang et al. 2023) which employ both the normal samples and a few anomaly data. Although these methods perform well in anomaly detection, they have limited performance in anomaly localization and can not deal with the task of anomaly classification (Hu et al. 2024). Therefore, researchers proposed anomaly generation methods to generate more anomaly data, to help achieve better performance by using supervised anomaly inspection.
Figure 1: Top: Evaluating anomaly generation quality in four aspects: Whether generate valid anomaly, degree of realism, alignment of the mask, and whether the location of mask is reasonable, the results show that our generated results are better than the other methods. Bottom: Our model can simultaneously generate extensive anomaly image-mask pairs.
Existing anomaly generation methods can be categorized into two groups, 1) model-free methods randomly cut and paste patches from existing anomalies or anomalous texture datasets onto normal samples (Lin et al. 2021; Li et al. 2021). But the anomaly data synthesized by them is unrealistic. 2) Generative methods employ generative models like GANs and diffusion models to generate anomaly data. Generally, The GAN-based model (Zhang et al. 2021; Niu et al. 2020) requires a large amount of training data to achieve better generation performance and they can not generate masks. DFMGAN (Duan et al. 2023) firstly trained on normal data and then migrated to anomaly data to achieve a few shot generation. This method also encounters the problem that the generated anomalies are not realistic enough and the masks are not sufficiently aligned because there is no explicit alignment constraint design. AnomalyDiffusion (Hu et al. 2024) based on the texture-inversion (Gal et al. 2022) technique of Diffusion (Rombach et al. 2022), separately learns the anomaly appearance and location information, then generates the anomaly on the masked normal samples. As AnomalyDiffusion only focuses on the part of anomaly, the generated anomalies do not blend realistically with the original image, and the masks generated individually may appear in the background of image.
To address these limitations, we propose DualAnoDiff, a novel few-shot anomaly image generation model that utilizes a dual-diffusion to simultaneously generate the overall image and the corresponding anomaly part. This approach can realize the effective integration of anomaly image and anomaly part, resulting in the generation of authentic and highly aligned anomaly image-mask data pairs. Specifically, our model is built upon a pre-trained diffusion model and introduces two LoRA to expand a single diffusion model into dual-interrelated diffusion model. One branch, referred to as the global branch, is responsible for generating the overall anomaly image, while the anomalous branch generates the localized anomaly image. They exchange information through the self-attention interaction module, This module combines the attention layers of global and anomalous branches and performs shared attention calculations, enabling the interaction and fusion of information in the dual-denoising process. This ensures the consistency between the generated overall anomaly image and the localized anomaly image. Furthermore, to further preserve the invariance of the background, we introduce a background compensation module based on self-attention adaptive injection. This module involves adding noise to the background image, extracting the key and value from the intermediate feature layer, and applying adaptive fusion MLP to incorporate the background information into the global branch. It helps to maintain consistency in the generated image and avoids coupling between the anomaly part and the background of the object in the image.
Fig. 1 shows our generation results outperform better compared with other method in four key aspects. Moreover, extensive experiments have been conducted in Mvtec AD (Bergmann et al. 2019) to quantitatively validate the superiority of the anomaly data generated by DualAnoDiff in downstream anomaly inspection tasks, and achieving a state-of-the-art level of performance with pixel-level 99.1% AUROC and 84.5% AP score.
Figure 2: The architecture of DualAnoDiff. 1) Two branches of DualAnoDiff generate the anomaly image and corresponding anomaly part simultaneously with different but nested prompts. 2) Two branches share the attention information after every attention block by Self-Attention Interaction Module (SAIM) during the denoising process to keep the consistency of generated images. 3) Background Compensation Module (BCM) extracts the Key, and Value of the background image and applies an adaptive fusion to SD, to help the model more focus on the object of the image.
Our contributions can be summarized as follows:
  • We propose DualAnoDiff, a novel few-shot diffusion-based anomaly generation method, which simultaneously generates both the overall image and the corresponding anomaly part with a highly aligned mask by a dual-interrelated diffusion model.
  • We design a background compensation approach, which involves image backgrounds as control information and injects the intermediate feature into the denoising process of anomaly image, to enhance the stability and authenticity of the generated data.
  • Extensive experiments demonstrate the superiority of our method over existing anomaly generation models in terms of both generation quality and performance of downstream anomaly inspection tasks.
Related Work
Few-shot Image Generation
Few-shot image generation aims to generate new and diverse examples while preventing overfitting to the few training images (Zhu et al. 2024). It is highly susceptible to overfit with extremely limited training data (less than 10) and then generate highly similar images. Early method (Mo, Cho, and Shin 2020) proposes modifying network weights, using various regularization techniques and data augmentation to prevent overfitting (Hu et al. 2024). There are also methods (Hu et al. 2023) to mitigate overfitting by pre-training on the source domain and subsequently migrating to fewer samples through cross-domain consistency losses to keep the generated distribution. Textual Inversion (Gal et al. 2022) and Dreambooth (Ruiz et al. 2023) encode a few images into the textual space of a pre-trained diffusion model to achieve diverse target customization generation which preserves its key visual features.
Anomaly Inspection
The anomaly inspection task consists of anomaly detection, localization, and classification (Duan et al. 2023). Due to the scarcity of abnormal data in industrial scenarios, most methods (Roth et al. 2022; Li et al. 2024) use unsupervised methods and semi-supervised methods. Reconstruction-based methods (Schlegl et al. 2019; Cao et al. 2023) detect anomalies by analyzing the residual image before and after reconstruction. Embedding-based methods (Lee, Lee, and Song 2022; Cao et al. 2022) utilize pre-trained networks to extract the image-level features and patch-level features, and then perform clustering according to the similarity between the features to detect the anomalies. All of those methods can only address the task of anomaly detection, while having limited performance in anomaly localization and being incapable of anomaly classification.
Anomaly Generation
Due to the scarcity of anomaly data, anomaly generation has emerged as a field of paramount importance. DRAEM (Zavrtanik, Kristan, and Skočaj 2021), Cut-Paste (Li et al. 2021), Crop-Paste (Lin et al. 2021) and PRN (Zhang et al. 2023) crop and paste unrelated textures or existing anomalies into normal sample. These approaches can be somewhat effective, but the generated anomalies are completely unrealistic and there is no way to use them as anomaly classification tasks. Subsequently, GANs (Goodfellow et al. 2020) have been applied for anomaly generation due to their ability to generate high-fidelity images. SDGAN (Niu et al. 2020) and DefectGAN (Zhang et al. 2021) generate anomalies on normal samples by learning from anomaly data. However, they require a large amount of anomaly data and cannot generate anomaly masks. DFMGAN (Duan et al. 2023) transfers a StyleGAN2 (Karras et al. 2020) pretained on normal samples to the anomaly domain, but lacks generation authenticity and accurate alignment between generated anomalies and masks. Subsequently, Diffusion models have been more widely used due to their extreme generalizability. AnomalyDiffusion (Hu et al. 2024) learns the features of anomaly and the distribution of masks through text inversion (Gal et al. 2022) technique of diffusion, to generate the specified anomaly at the position of the corresponding mask of the normal image. However, Since the method learns the anomaly part and the mask separately, it makes the generated mask not necessarily located on the object, and the anomaly-object transition is not natural enough. Our approach is similarly based on the diffusion model but achieves more realistic, distinct, and diverse anomaly data generation.
Method
Given a set of a limited number of anomaly image-mask pairs, our goal is to learn the features of the anomaly images, and then generate more anomaly image-mask pairs that belong to the same item and anomaly type, while ensuring greater diversity in the distribution and appearance of the anomaly. In our method, we generate the overall image and the anomaly part, which is then segmented to get the corresponding mask.
Preliminaries
Latent Diffusion Models. Stable Diffusion (SD), a variant of the latent diffusion model (LDM) (Rombach et al. 2022), serves as a text-guided diffusion model. To generate high-resolution images while enhancing computational efficiency in the training process, it employs a pre-trained variational autoencoder (VAE) (Kingma and Welling 2013) encoder $\mathcal{E}(\cdot)$ to map images into latent space and perform an iterative denoising process. Subsequently, the predicted images are mapped back into pixel space through the pre-trained VAE decoder $D(\cdot)$. $\epsilon_{\theta}$ is the denoising network, for each denoising step, the simplified optimization objective is defined as follows:
$L_{LDM}(\theta)=\mathbb{E}_{\mathcal{E}(x),\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}\left(z_{t},t,\tau_{\theta}(c)\right)\|_{2}^{2}\right]$ (1)
where $\epsilon$ are latent noise, the text description $c$ is encoded by the CLIP (Radford et al. 2021) text encoder $\tau_{\theta}(\cdot)$ and then used to guide the diffusion denoising process.
DualAnoDiff Framework
We have been searching for methods that can generate both anomaly image and mask, inspired by Layerdiffusion (Zhang and Agrawala 2024), we decompose an anomaly image into two parts, the overall anomaly image and the corresponding anomaly part, where the overall image refers to the whole anomaly image $I$, the anomaly part refers to the part that contains only the anomaly $I_{a}$. $I_{a}=I\times M_{a}$ ($M_{a}$ is the mask of anomaly).
As shown in Fig. 2, the proposed DualAnoDiff involves two interrelated diffusion models, and they share part information by the Self-attention Interaction Module. We freeze the weights of diffusion models and use two LORA (Hu et al. 2021) to fine-tune them. For ease of description, we denote the two diffusion models as $SD$ and $SD^{\ast}$, $SD$ denote the diffusion model to generate image $I$, $SD^{\ast}$ denote the diffusion model to generate $I_{a}$.
Dual-Interrelated Diffusion. AnomalyDiffusion (Hu et al. 2024) primarily focuses on the anomaly part, which may result in generated anomaly images lacking a convincingly realistic appearance. Directly generating the complete anomaly image poses challenges in obtaining the corresponding mask. To address those limitations, our proposed model simultaneously generates both the overall image and the anomaly part. This novel approach overcomes challenges in generating realistic anomaly images while ensuring the availability of accurate masks.
Firstly, we encode $I$, $I_{a}$ into latent space $z$,$z^{\prime}$ = $\varepsilon(I),\varepsilon(I_{a})$ by using VAE encoder $\varepsilon(\cdot)$, employ a forward process to add noise into the latents with the same timestep $t$, and then learn to denoising during the backward process guided by different prompts. Throughout these processes, information is shared and synchronized between the two diffusion models through the SAIM, enabling the model to effectively fit the training data pairs. By generating the anomalies separately, this approach achieves two important objectives. Firstly, it enhances the diversity and reality of the generated anomaly through a simple yet effective operation of adding two LoRA. Secondly, it ensures a highly aligned mask that accurately corresponds to the anomaly image.
Nested Prompts. The goal of the dual diffusion is to generate an anomaly image and anomaly part pair ($I$ and $I^{\prime}$), which exhibits an inclusion relationship. To facilitate the model’s understanding of the distinct entities within the image (the primary subject and the anomaly), we employ a pair of prompts designed to reflect this inclusion relationship:
$p:\medspace$ $a\medspace\mathbf{x}\medspace with\medspace\mathbf{y}$
$p^{\prime}:\medspace$ $\mathbf{y}$ (2)
where the prompts $p$ and $p^{\prime}$ correspond to the anomaly image $I$ and the corresponding anomaly part $I_{a}$ respectively. Both prompts are encoded by the text encoder $\tau_{\theta}(\cdot)$ and then injected into Unet (Ronneberger, Fischer, and Brox 2015) of Diffusion. $\tau_{\theta}(\cdot)$ is trainable.
Figure 3: b, c are the cross attention maps of different text tokens in SD corresponding to the text of “a vfx with” and “sks”. a, d are the generated image from SD and SD*.
The variables $x$ and $y$ can be the class name and anomaly name provided by the dataset. In our model, we use the $vfx$ and $sks$ which were suggested by DreamBooth (Ruiz et al. 2023). Those words have weak prior in both the language model and the diffusion model, making them easier to fit than other words, and can achieve better generation results, specially for high-prior words. Fig. 3 presents the generated result and visualizes the cross-attention maps between text token and vision. where $a$, $d$ are the anomaly image and anomaly part generated by $SD$ and $SD^{\ast}$. $b$ and $c$ are the 64 $\times$ 64 resolutions feature maps randomly extracted from the second half of the generation process in the U-Net of $SD$. Where $b$ corresponds to the text “$a$ $vfx$ $with$”, $c$ corresponds to “$sks$”. It is evident that the model correctly separates the attributes of anomaly and object, and accurately associating them to the specified text as we want.
Category CDC Crop-Paste SDGAN Defect-GAN DFMGAN AnomalyDiffusion Ours
IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$ IS $\uparrow$ IC-L $\uparrow$
bottle 1.52 0.04 1.43 0.04 1.57 0.06 1.39 0.07 1.62 0.12 1.58 0.19 2.17 0.36
cable 1.97 0.19 1.74 0.25 1.89 0.19 1.70 0.22 1.96 0.25 2.13 0.41 2.12 0.43
capsule 1.37 0.06 1.23 0.05 1.49 0.03 1.59 0.04 1.59 0.11 1.59 0.21 1.60 0.31
carpet 1.25 0.03 1.17 0.11 1.18 0.11 1.24 0.12 1.23 0.13 1.16 0.24 1.36 0.29
grid 1.97 0.07 2.00 0.12 1.95 0.10 2.01 0.12 1.97 0.13 2.04 0.44 2.09 0.42
hazel_nut 1.97 0.05 1.74 0.21 1.85 0.16 1.87 0.19 1.93 0.24 2.13 0.31 1.91 0.35
leather 1.80 0.07 1.47 0.14 2.04 0.12 2.12 0.14 2.06 0.17 1.94 0.41 1.88 0.34
metal_nut 1.55 0.04 1.56 0.15 1.45 0.28 1.47 0.30 1.49 0.32 1.96 0.30 1.56 0.32
pill 1.56 0.06 1.49 0.11 1.61 0.07 1.61 0.10 1.63 0.16 1.61 0.26 1.82 0.37
screw 1.13 0.11 1.12 0.16 1.17 0.10 1.19 0.12 1.12 0.14 1.28 0.30 1.34 0.36
tile 2.10 0.12 1.83 0.20 2.53 0.21 2.35 0.22 2.39 0.22 2.54 0.55 2.35 0.50
toothbrush 1.63 0.06 1.30 0.08 1.78 0.03 1.85 0.03 1.82 0.18 1.68 0.21 2.40 0.48
transistor 1.61 0.13 1.39 0.15 1.76 0.13 1.47 0.13 1.64 0.25 1.57 0.34 1.69 0.33
wood 2.05 0.03 1.95 0.23 2.12 0.25 2.19 0.29 2.12 0.35 2.33 0.37 2.21 0.40
zipper 1.30 0.05 1.23 0.11 1.25 0.10 1.25 0.10 1.29 $\underline{0.27}$ 1.39 0.25 2.09 0.36
Average 1.65 0.07 1.51 0.14 1.71 0.13 1.69 0.15 1.72 0.20 1.80 0.32 1.90 0.37
Table 1: Comparison on IS and IC-LPIPS on MVTec dataset. Our model generates the most high-quality and diverse anomaly data, achieving the best IS and IC-LPIPS. Bold and underline represent optimal and sub-optimal results, respectively.
Self-attention Interaction Module (SAIM). During training, $SD$ and $SD^{\ast}$ share the same timestep $t$ and denoising simultaneously, they share information after every attention in the unet. For example, self-attention is more likely to share the positional information and cross-attention shares the semantic information. The shared step is formulated as:
$\varphi_{i}(\tilde{z})=$ $Rearrange(Cancat(\varphi_{i}(z),\varphi_{i}(z^{\prime})))$
$\varphi_{i}(\tilde{z})_{new}=$ $SelfAtt(\varphi_{i}(\tilde{z}))$
$\varphi_{i}(z),\varphi_{i}(z^{\prime})=$ $Split(Rearrange(\varphi_{i}(\tilde{z})_{new}+\varphi_{i}(\tilde{z})))$ (3)
where $\varphi^{i}$ is the intermediate representation of the U-Net. The original shape of $\varphi_{i}(z)$ and $\varphi_{i}(z^{\prime})$ are “b w c”, where b denotes the batch size, w represents the spatial dimension, and c represents the channel dimension. We rearrange the $cancat(\varphi_{i}(z),\varphi_{i}(z^{\prime}))$ to shape of “bw 2 c”. Concat and split are a set of corresponding operations that are used to aggregate and separate $varphi_{i}(z),\varphi_{i}(z^{\prime}))$.
Figure 4: Comparison between the models without (a) and with (b) the Background Compensation Module
Loss Function. With the two-stream parallel structure of diffusion, our final training objective expressed as follows:
$\mathcal{L}=$ $\mathbb{E}_{\mathcal{E}(I),\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}\left(z_{t},t,\tau_{\theta}(p))\right)\|_{2}^{2}\right]$
$+\mathbb{E}_{\mathcal{E}(I_{a}),\epsilon^{\ast},t}\left[\|\epsilon^{\ast}-\epsilon_{\theta}^{\ast}\left(z^{\prime}_{t},t,\tau_{\theta}(p^{\prime})\right)\|_{2}^{2}\right]$ (4)
here, $\epsilon$ and $\epsilon^{\ast}$ are latent noise for the anomaly image and anomaly part, $t$ is the time step, $p$, $p^{\prime}$ are the corresponding prompts, they are encoded by text encoder $\tau_{\theta}(\cdot)$ which is trainable.
Mask Generation. Thanks to our two-stream parallel structure, which generates the anomaly part as a single entity, obtaining the precise mask becomes straightforward. There are two ways to obtain a high-quality mask. Firstly, we can utilize existing segmentation algorithms such as SAM (Kirillov et al. 2023), U2-Net (Qin et al. 2020) to obtain high-quality masks after generating the anomaly part image. Additionally, during the generation process, we can extract the average attention maps in $SD^{\ast}$ to compute the mask such as (Wang et al. 2023) which is widely used in semantic segmentation. In our paper, we use the first method.
Background Compensation Module
Although the model has achieved good results so far, there are still some problems in some cases, which is caused by the limited training data. Several bad cases of corresponding generated results are shown in Fig. 4 for categories bottle, pill, and toothbrush in the Mvtec (the average training data for each category is 8). The bottle was only partially generated. The edges of pill is blurred and the internal properties leakage into the background. Additionally, the toothbrush has two brush handles, which is an abnormal occurrence. Furthermore, all of those generated images lack sharpness and the background color does not maintain sufficient purity. In general, the model is incapable of fully grasping the characteristics of those cases, there are problems such as objects and background mixed, and objects deformed.
To enhance the model’s ability to learn the shape of object and focus more on object generation, we design a background compensation module. First, We employ U2-Net to segment the image $I$ and obtain the object mask $M_{f}$, then get the background image $I_{b}=(1-M_{f})\times I$, $I_{b}$ will process the $SD$ as same as $I$ (except for the SAIM). Next, We utilize $I_{b}$ as a condition, which contains both background and mask information. This allows us to to control the shape of the object while providing background context. We collect its features in self-attentions and inject them into the $K$, $V$ of the $I$ in every corresponding self-attention steps. The injection process can be formulated as:
$\varphi_{i}(z^{b})=$ $SelfAtt(\varphi_{i}(z^{b}))$
$Q=$ $W^{(i)}_{Q}\cdot\varphi_{i}(z)$
$\varphi_{i}(z)=$ $\varphi_{i}(z)+\gamma MLP(\varphi_{i}(z^{b}))$
$K=W^{(i)}_{Q}\cdot$ $\varphi_{i}(z)\quad V=W^{(i)}_{Q}\cdot\varphi_{i}(z)$ (5)
where $\gamma$ is a learnable scale factor initialized to be 0.1. This design can help us to maximally preserve the generative effect of the mainstream SD, meanwhile use background information and the shape of the mask. With the condition of $I_{b}$, the loss function becomes:
$\mathcal{L}=$ $\mathbb{E}_{\mathcal{E}(I),\epsilon,t}\left[\left\|\epsilon-\epsilon_{\theta}\left(z_{t},t,\tau_{\theta}(p),I_{b}\right)\right\|_{2}^{2}\right]$
$+\mathbb{E}_{\mathcal{E}(I),\epsilon^{\ast},t}\left[\left\|\epsilon^{\ast}-\epsilon_{\theta}^{\ast}\left(z^{\prime}_{t},t,\tau_{\theta}(p^{\prime})\right)\right\|_{2}^{2}\right]$ (6)
Category DRAEM PRN DFMGAN AnomalyDiffusion Ours
AUC-P AP-P $F_{1}$-P AP-I AUC-P AP-P $F_{1}$-P AP-I AUC-P AP-P $F_{1}$-P AP-I AUC-P AP-P $F_{1}$-P AP-I AUC-P AP-P $F_{1}$-P AP-I
bottle 96.7 80.2 74.0 99.8 97.5 76.4 71.3 98.4 98.9 90.2 83.9 99.8 99.4 94.1 87.3 99.9 99.5 93.4 85.7 100
cable 80.3 21.8 28.3 83.2 94.5 64.4 61.0 92.0 97.2 81.0 75.4 97.8 99.2 90.8 83.5 100 97.5 82.6 76.9 98.3
capsule 76.2 25.5 32.1 98.7 95.6 45.7 47.9 95.8 79.2 26.0 35.0 98.5 98.8 57.2 59.8 99.9 99.5 73.2 67.0 99.2
carpet 92.6 43.0 41.9 98.7 96.4 69.6 65.6 97.8 90.6 33.4 38.1 98.5 98.6 81.2 74.6 98.8 99.4 89.1 80.2 99.9
grid 99.1 59.3 58.7 99.9 98.9 58.6 58.9 98.9 75.2 14.3 20.5 90.4 98.3 52.9 54.6 99.5 98.5 57.2 54.9 99.7
hazelnut 98.8 73.6 68.5 100 98.0 73.9 68.2 96.0 99.7 95.2 89.5 100 99.8 96.5 90.6 99.9 99.8 97.7 92.8 100
leather 98.5 67.6 65.0 100 99.4 58.1 54.0 99.7 98.5 68.7 66.7 100 99.8 79.6 71.0 100 99.9 88.8 78.8 100
metal_nut 96.9 84.2 74.5 99.6 97.9 93.0 87.1 99.5 99.3 98.1 94.5 99.8 99.8 98.7 94.0 100 99.6 98.0 93.0 99.9
pill 95.8 45.3 53.0 98.9 98.3 55.5 72.6 97.8 81.2 67.8 72.6 91.7 99.8 97.0 90.8 99.6 99.6 95.8 89.2 99.0
screw 91.0 30.1 35.7 96.3 94.0 47.7 49.8 94.7 58.8 2.2 5.3 64.7 97.0 51.8 50.9 97.9 98.1 57.1 56.1 95.0
tile 98.5 93.2 87.8 100 98.5 91.8 84.4 96.9 99.5 97.1 91.6 100 99.2 93.9 86.2 100 99.7 97.1 91.0 100
toothbrush 93.8 29.5 28.4 99.8 96.1 46.4 46.2 100 96.4 75.9 72.6 100 99.2 76.5 73.4 100 98.2 68.3 68.6 99.7
transistor 76.5 31.7 24.2 80.5 94.9 68.6 68.4 88.9 96.2 81.2 77.0 92.5 99.3 92.6 85.7 100 98.0 86.7 79.6 93.7
wood 98.8 87.8 80.9 100 96.2 74.2 67.4 92.7 95.3 70.7 65.8 99.4 98.9 84.6 74.5 99.4 99.4 91.6 83.8 99.9
zipper 93.4 65.4 64.7 100 98.4 79.0 73.7 99.7 92.9 65.6 64.9 99.9 99.4 86.0 79.2 100 99.6 90.7 82.7 100
Average 92.2 54.1 53.1 97.0 96.9 66.2 64.7 96.6 90.0 62.7 62.1 94.8 99.1 81.4 76.3 99.7 99.1 84.5 78.8 98.9
Table 2: Comparison on pixel-level and image-level anomaly localization on MVTec dataset by training an U-Net on the generated data from DRAEM, PRN, DFMGAN, AnomalyDiffusion and our model.
Experiments
Experiment Settings
Dataset. we conduct experiments on MVTec (Bergmann et al. 2019). We refer to the experimental setup of AnomalyDiffusion (Hu et al. 2024) using 1/3 of the data as the training set, reserving the remaining two-thirds for testing. Implementation details. We train a model for each anomaly type separately and generate 500 anomaly image-mask pairs for the downstream anomaly inspection tasks. More details of the experiment can be found in the supplementary materials. Metric. 1) For generation, we use Inception Score (IS) and Intra-cluster pairwise LPIPS distance (IC-LPIPS) (Ojha et al. 2021) as AnomalyDiffusion (Hu et al. 2024) to measure the generation quality and generation diversity. 2) For anomaly inspection. We use AUROC, Average Precision (AP), and the F1-max score to evaluate the accuracy of anomaly detection and localization.
Comparison in Anomaly Generation
Baseline. 1) We select DRAEM(Zavrtanik, Kristan, and Skočaj 2021), PRN (Zhang et al. 2023), DFMGAN (Duan et al. 2023) and AnomalyDiffusion (Hu et al. 2024) which can generate anomaly image-mask pairs as comparative benchmarks to compare anomaly detection and localization. 2) We choose CDC (Ojha et al. 2021), Crop-Paste, SDGAN (Niu et al. 2020), DefectGAN (Zhang et al. 2021), DFMGAN (Duan et al. 2023) and AnomalyDiffusion (Hu et al. 2024) that can only generate specific anomaly types to compare anomaly generation quality and classification.
Figure 5: Comparison of the generation results on MVTec. Our model excels in generating high-quality anomaly images that are accurately aligned with the anomaly masks.
Category DiffAug Crop-Paste DFMGAN AnoDiff Ours
bottle 48.84 52.71 56.59 90.70 76.74
cable 21.36 32.81 45.31 67.19 71.88
capsule 34.67 32.8 37.23 66.67 70.67
carpet 35.48 27.96 47.31 58.06 74.19
grid 28.33 28.33 40.83 42.50 80.0
hazelnut 65.28 59.03 81.94 85.42 89.58
leather 40.74 34.39 49.73 61.90 90.48
metalnut 58.85 59.89 64.58 59.38 82.81
pill 29.86 26.74 29.52 59.38 54.17
screw 25.10 28.81 37.45 48.15 70.37
tile 59.65 68.42 74.85 84.21 100
transistor 38.09 41.67 52.38 60.71 67.86
wood 41.27 47.62 49.21 71.43 85.71
zipper 22.76 26.42 27.64 69.51 69.51
Average 39.31 40.55 49.61 66.09 72.26
Table 3: Comparison on anomaly classification accuracy trained on the generated data by the anomaly generation models with a ResNet-18.
Anomaly Generation Quality. Tab. 1 shows the results of the quantification metrics for image quality. For each anomaly category, we allocate one-third of the anomaly data for training and generate 500 anomaly images to compute IS and IC-LPIPS. It demonstrates that our model generates anomaly data with both the highest quality and diversity.
Category Unsupervised Supervised
DRAEM SSPCAB CFA RD4AD PatchCore Musc DevNet DRA PRN Ours
bottle 99.1/88.5 98.9/88.6 98.9/50.9 98.8/51.0 97.6/75.0 98.5/82.8 96.7/67.9 91.7/41.5 99.4/92.3 99.5/93.4
cable 94.8/61.4 93.1/52.1 98.4/79.8 98.8/77.0 96.8/65.9 96.2/58.8 97.9/67.6 86.1/34.8 98.8/78.9 97.5/82.6
capsule 97.6/47.9 90.4/48.7 98.9/71.1 99.0/60.5 98.6/46.6 98.9/52.7 91.1/46.6 88.5/11.0 98.5/62.2 99.5/73.2
carpet 96.3/62.5 92.3/49.1 99.1/47.7 99.4/46.0 98.7/65.0 99.4/75.3 94.6/19.6 98.2/54.0 99.0/82.0 99.4/89.1
grid 99.5/53.2 99.6/58.2 98.6/82.9 98.0/75.4 97.2/23.6 98.6/37.0 90.2/44.9 86.2/28.6 98.4/45.7 98.5/57.2
hazelnut 99.5/88.1 99.6/94.5 98.5/80.2 94.2/57.2 97.6/55.2 99.3/74.4 76.9/46.8 88.8/20.3 99.7/93.8 99.8/97.7
leather 98.8/68.5 97.2/60.3 96.2/60.9 96.6/53.5 98.9/43.4 99.7/62.6 94.3/66.2 97.2/ 5.1 99.7/69.7 99.9/88.8
metal_nut 98.7/91.6 99.3/95.1 98.6/74.6 97.3/53.8 97.5/86.8 87.5/49.6 93.3/57.4 80.3/30.6 99.7/98.0 99.6/98.0
pill 97.7/44.8 96.5/48.1 98.8/67.9 98.4/58.1 97.0/75.9 97.6/65.6 98.9/79.9 79.6/22.1 99.5/91.3 99.6/95.8
screw 99.7/72.9 99.1/62.0 98.7/61.4 99.1/51.8 98.7/34.2 98.9/31.7 66.5/21.1 51.0/5.1 97.5/44.9 98.1/57.1
tile 99.4/96.4 99.2/96.3 98.6/92.6 97.4/78.2 94.9/56.0 98.3/80.6 88.7/63.9 91.0/54.4 99.6/96.5 99.7/97.1
toothbrush 97.3/49.2 97.5/38.9 98.4/61.7 99.0/63.1 97.6/37.1 99.4/64.2 96.3/52.4 74.5/4.8 99.6/78.1 98.2/68.3
transistor 92.2/56.0 85.3/36.5 98.6/82.9 99.6/50.3 91.8/66.7 92.5/61.2 55.2/4.4 79.3/11.2 98.4/85.6 98.0/86.7
wood 97.6/81.6 97.2/77.1 97.6/25.6 99.3/39.1 95.7/54.3 98.7/77.5 93.1/47.9 82.9/21.0 97.8/82.6 99.4/91.6
zipper 98.6/73.6 98.1/78.2 95.9/53.9 99.7/52.7 98.5/63.1 98.4/64.2 92.4/53.1 96.8/42.3 98.8/77.6 99.6/90.7
Average 97.7/69.0 96.2/65.5 98.3/66.3 98.3/57.8 97.1/56.6 97.5/62.6 86.4/49.3 84.8/25.7 99.0/78.6 99.1/84.5
Table 4: Comparison on pixel-level anomaly localization (AUROC/AP) between the simple U-Net trained on our generated dataset and the existing anomaly detection methods with their official codes or pre-trained models.
Moreover, we present the anomaly images generated by several prominent anomaly image generation models in Fig. 5, among which, AnomalyDiffusion is by far the stat-of-art method. It is observed that the images generated by Defect-GAN often do not have anomalies, DFMGAN tends to introduce noise in the image, and the anomaly image-mask pairs lack proper alignment. The state-of-the-art model AnomalyDiffusion may generates an unreasonable mask that is outside the object, and the generated anomaly does not fit the original image enough to make the anomaly look unreal. In contrast, our proposed model demonstrates the ability to generate highly realistic and valid anomalies, even excelling in handling smaller-scale anomalies such as grid-bent. More results are presented in the supplementary material.
Anomaly Generation for Anomaly Detection and Localization. we evaluate the effectiveness of our approach by comparing it with existing methods for anomaly generation in downstream anomaly detection and localization. We calculate pixel-level dimensions and image-level AUROC, AP, and F1-max for each category. The results are presented in Tab. 2, and because the image-level results of AUROC, AP, and F1-max are relatively similar, as well as the limitation of the table width, we only show the results of the most representative AP, and the other results of AUROC and F1-max are shown in the supplementary material. It can be seen that the U-net trained in our generated data reaches the highest AP of 84.5% and F1-max of 78.8%, surpassing the second-ranked AnomalyDiffusion by a margin of 3.1%(AP).
Anomaly Generation for Anomaly Classification. To further assess the quality of the anomaly images generated by our model, we utilize them to train a downstream anomaly classification model. Following the experimental setup of DFMGAN (Duan et al. 2023), we employ ResNet-34 to train on the generated dataset. We then evaluate the classification accuracy on our test dataset as shown in Tab. 3. As evident from the data, our model demonstrates significantly higher accuracy across the majority of categories compared to other models, with an average accuracy improvement of 6.17%, indicating the anomaly data we generate is more realistic.
Figure 6: Comparison of generation results for different prompts on the hazelnut-print category.
Comparison with Anomaly Detection Models
To further validate the effectiveness of our model, we compare it with state-of-the-art anomaly detection methods DRAEM(Zavrtanik, Kristan, and Skočaj 2021), SSPCAB (Ristea et al. 2022), CFA (Lee, Lee, and Song 2022), RD4AD (Deng and Li 2022), PatchCore (Roth et al. 2022), MuSc (Li et al. 2024), DevNet (Pang et al. 2021), DRA (Ding, Pang, and Shen 2022) and PRN (Zhang et al. 2023). We utilize their official codes or pre-trained models and evaluate them on the same testing dataset (2/3 anomaly data in the test of Mvtec). As there is no available open-source code for PRN, we rely on the data provided in its research paper. The methods are evaluated based on their pixel-level AUROC and AP scores, as demonstrated in Tab. 4.
Category Without BCM With BCM
AUROC AP $F_{1}$-max AUROC AP $F_{1}$-max
bottle 98.4 88.8 77.1 99.5 93.4 85.7
pill 98.4 86.9 78.2 99.6 95.8 89.2
toothbrush 97.2 62.7 64.0 98.2 68.3 68.6
Average 98.8 83.0 78.1 99.1 84.5 78.8
Table 5: Comparison on pixel-level anomaly localization on part of categories in MVTec between with/without the Background Compensation Module. Average is the mean of all categories in Mvtec.
Ablation Study
Fig. 6 presents the results of using different prompts in our model. It is evident that using the prompt “a vfx with sks” yields superior performance in terms of object shape, color, and other aspects compared to using the category name directly. More quantitative indicators can be found in the supplementary material. Furthermore, we assess the effectiveness of the Background Compensation Module (BCM), which exhibits more noticeable improvements in the categories of bottle, pill, and toothbrush. The visualization results are depicted in Fig. 4, and the pixel-level localization results are presented in Tab. 5.
Conclusion
In this paper, we present a novel approach, DualAnoDiff, for generating anomalous image-mask pairs. Our method employs a parallel dual-diffusion to simultaneously generate the anomaly image and the corresponding anomaly part image. This ensures a high level of alignment between the generated anomaly image-mask pair and the authenticity of the anomaly. Additionally, to address challenging cases, we introduce a background compensation module that efficiently enhances the model’s fitting capability. Extensive experiments demonstrate its superior performance compared to existing anomaly generation methods. The anomaly data generated by our model effectively enhances the performance of downstream anomaly inspection tasks.